OpenStudy (shelby1290):

**** WILL FAN AND MEDAL **** Determine whether each quadratic function intersects the x-axis at one point, two points or not at all. a) g(x)= 4(x-6)^2+2 b) f(x)=-3(x-5)^2+2 c) f(x)= -3(x+2)^2

OpenStudy (anonymous):

It going to take a while to the answer so please wait patiently :3

OpenStudy (shelby1290):

@TypnoEchosis Do you understand how to solve these questions?

OpenStudy (anonymous):

Yes O 3O

OpenStudy (shelby1290):

well for part a) I'm not sure if i have to foil or not

OpenStudy (anonymous):

Oh okay

OpenStudy (shelby1290):

So, do i have to use the foil method?

OpenStudy (anonymous):

Yes use the foil method

OpenStudy (anonymous):

Do both the foil and non-foil method and see what you get. It really helps :)

OpenStudy (shelby1290):

g(x)=4(x-6)^2+2 =4(x-6)(x-6)+2 =4(x^2-6x-6x+36)+2 =4(x^2-12x+36)+2 What do i do next? do i distribute 4 into the brackets? Or factor? ( finding two numbers which multiply to 36 and add to -12)

OpenStudy (shelby1290):

@TypnoEchosis ^^

OpenStudy (anonymous):

Yes distrubute 4 into brackets

OpenStudy (anonymous):

Doing this is basically like doing order of operations except much more complicated.

OpenStudy (shelby1290):

=4(x^2-12x+36)+2 =4x^2-48x+144+2 =4x^2-48x+146

OpenStudy (shelby1290):

b^2-4ac=(-48)^2-4(4)(146) =2,304-2,336 =-32 @TypnoEchosis The parabola will not intersect the x-axis?

OpenStudy (shelby1290):

is that right?

OpenStudy (anonymous):

a) not at all b) two point intersection c) one point (tangent to the x axis at one point) A two page solution with plots is attached.

OpenStudy (anonymous):

The attachment was produced by Mathematica v9 .

OpenStudy (shelby1290):

thank you!